# Low-Dimensional Execution Manifolds in Transformer Learning Dynamics: Evidence from Modular Arithmetic Tasks

### v1

@@ Abstract
We investigate the geometric structure of learning dynamics in overparameterized transformer models through carefully controlled modular arithmetic tasks. Our primary finding is that despite operating in high-dimensional parameter spaces ($d=128$), transformer training trajectories rapidly collapse onto low-dimensional execution manifolds of dimension 3344. This dimensional collapse is robust across random seeds and moderate task difficulties, though the orientation of the manifold in parameter space varies between runs. We demonstrate that this geometric structure underlies several empirically observed phenomena: (1) sharp attention concentration emerges as saturation along routing coordinates within the execution manifold, (2) stochastic gradient descent (SGD) exhibits approximately integrable dynamics when projected onto the execution subspace, with non-integrability confined to orthogonal staging directions, and (3) sparse autoencoders capture auxiliary routing structure but fail to isolate execution itself, which remains distributed across the low-dimensional manifold. Our results suggest a unifying geometric framework for understanding transformer learning, where the vast majority of parameters serve to absorb optimization interference while core computation occurs in a dramatically reduced subspace. These findings have implications for interpretability, training curriculum design, and understanding the role of overparameterization in neural network learning.
@@ 1.1 Contributions
  • Localization of SGD non-integrability: SGD commutators $\nabla_{A},\nabla_{B}]=\theta_{AB}-\theta_{BA}$ are large in ambient space but negligible when projected onto execution directions, indicating approximately integrable dynamics along the computational trajectory.
@@ 3.3.2 SGD Commutator Analysis
Projection decomposition. Let $P$ denote the projection operator onto the learned execution subspace (the top $k$ principal components from PCA). We decompose commutators:
$[\nabla_{A},\nabla_{B}]$ $=P([\nabla_{A},\nabla_{B}])+(I-P)([\nabla_{A},\nabla_{B}])$ (6)
@@
The ratio $\|[\nabla_{A},\nabla_{B}]_{\parallel}\|/\|[\nabla_{A},\nabla_{B}]\|$ quantifies what fraction of non-commutativity lives within the execution manifold.
@@ 4.3 Localization of SGD Non-Integrability
SGD updates in the full high-dimensional parameter space exhibit strong non-commutativity. Figure 3 shows raw commutator norms $\|[\nabla_{A},\nabla_{B}]\|$ throughout training. These norms remain large and frequently spike, even in later training stages when loss has largely plateaued. The spikes correlate only weakly with loss fluctuations, suggesting that non-commutativity is not simply a manifestation of optimization difficulty.
Projection reveals integrability. However, projecting these commutators onto the learned execution subspace reveals a dramatically different picture. Figure 4 shows the ratio of projected commutator norm to raw commutator norm:
$\rho=\frac{\|[\nabla_{A},\nabla_{B}]_{\parallel}\|}{\|[\nabla_{A},\nabla_{B}]\|}$ (10)
This ratio rapidly decreases and stabilizes near zero, indicating that when restricted to execution directions, SGD dynamics become approximately integrable. Nearly all non-commutativity lives in the orthogonal complement—the vast “staging” space perpendicular to the execution manifold.
Figure 3: Raw commutator norms $\|[\nabla_{A},\nabla_{B}]\|$ remain large throughout training, with frequent spikes indicating strong non-commutativity in the full parameter space.
Figure 4: The ratio $\rho=\|[\nabla_{A},\nabla_{B}]_{\parallel}\|/\|[\nabla_{A},\nabla_{B}]\|$ rapidly approaches zero, indicating that non-commutativity is confined to directions orthogonal to the execution manifold.
Explicit decomposition. Figure 5 presents an explicit decomposition of commutators into parallel and perpendicular components. The perpendicular component accounts for $95\%$ of total commutator magnitude throughout most of training. This localization of non-integrability has profound implications: the execution of the task proceeds via approximately commuting updates, while optimization interference is absorbed in orthogonal directions that don’t directly affect computation.
@@
This finding suggests a geometric role for overparameterization: extra dimensions provide a “buffer space” where optimization noise, task interference, and stochastic fluctuations can be sequestered without disrupting the core computational trajectory. The execution manifold acts as a stable attractor along which learning progresses in an orderly, approximately path-independent manner.
Figure 5: Decomposition of commutators into components parallel ($\parallel$) and perpendicular ($\perp$) to the execution manifold. The perpendicular component dominates, accounting for $95\%$ of total magnitude.
@@ 4.6 Extension to Modular Multiplication
Preliminary experiments with modular multiplication confirm that the execution manifold framework extends beyond addition. Multiplication tasks converge to slightly higher-dimensional manifolds (4466 dimensions versus 3344 for addition), with dimensionality scaling predictably with computational complexity. All attention matrices again operate in low-dimensional subspaces, and the same geometric principles apply: commutator null space structure, approximately integrable dynamics along execution directions, and curriculum effects. This suggests that manifold dimensionality may be a robust, task-dependent quantity that reflects inherent computational structure.
@@ 5.1 Theoretical Interpretation: Overparameterization and Parallel Exploration
The localization of non-integrability to orthogonal directions provides a mechanism for this stabilization: once the execution manifold is discovered, updates within it approximately commute, preventing destructive interference. Meanwhile, staging directions absorb the non-commutative aspects of SGD, allowing continued exploration and refinement without disrupting the core solution.
@@ 6 Conclusion
We have presented evidence that transformer learning dynamics collapse onto low-dimensional execution manifolds, with dimension 3344 for modular addition tasks despite ambient parameter space dimension $128$. This geometric structure unifies several empirical phenomena: attention bubbling emerges as saturation along routing coordinates, SGD exhibits approximately integrable dynamics when projected onto the execution subspace, and sparse features capture auxiliary routing structure distinct from distributed execution.
The localization of SGD non-commutativity to orthogonal staging directions suggests a fundamental role for overparameterization: extra dimensions absorb optimization interference, allowing core computation to proceed via approximately path-independent updates. Training curriculum critically determines whether models discover unified execution manifolds supporting compositional generalization or become trapped in fragmented solution regions.

### v2

@@ Abstract
We investigate the geometric structure of learning dynamics in overparameterized transformer models through carefully controlled modular arithmetic tasks. Our primary finding is that despite operating in high-dimensional parameter spaces ($d=128$), transformer training trajectories rapidly collapse onto low-dimensional execution manifolds of dimension 3344. This dimensional collapse is robust across random seeds and moderate task difficulties, though the orientation of the manifold in parameter space varies between runs. We demonstrate that this geometric structure underlies several empirically observed phenomena: (1) sharp attention concentration emerges as saturation along routing coordinates within the execution manifold, (2) SGD commutators are preferentially aligned with the execution subspace (up to $10\times$ random baseline) early in training, with $92\%$ of non-commutativity confined to orthogonal staging directions and this alignment decreasing as training converges, and (3) sparse autoencoders capture auxiliary routing structure but fail to isolate execution itself, which remains distributed across the low-dimensional manifold. Our results suggest a unifying geometric framework for understanding transformer learning, where the vast majority of parameters serve to absorb optimization interference while core computation occurs in a dramatically reduced subspace. These findings have implications for interpretability, training curriculum design, and understanding the role of overparameterization in neural network learning.
@@ 1.1 Contributions
[…]
@@ 3.3.2 SGD Commutator Analysis
Projection decomposition. Let $B\in\mathbb{R}^{P\times K}$ be an orthonormal basis for the learned execution subspace, constructed from the top $K$ principal components of each attention block’s weight trajectory (Section 3.3.1), embedded in the full parameter space and QR-orthonormalized. The projection $\Pi=BB^{T}$ decomposes commutators:
$[\nabla_{A},\nabla_{B}]$ $=\Pi([\nabla_{A},\nabla_{B}])+(I-\Pi)([\nabla_{A},\nabla_{B}])$ (6)
@@
The projection fraction $\rho_{\mathrm{exec}}=\|[\nabla_{A},\nabla_{B}]_{\parallel}\|/\|[\nabla_{A},\nabla_{B}]\|$ quantifies what fraction of non-commutativity lives within the execution manifold.
Scale normalization. We report both the raw commutator $\delta=\theta_{AB}-\theta_{BA}$ and the normalized defect:
$D=\frac{\|\delta\|}{\|\eta\,\nabla_{\mathcal{L}}(\theta_{0};A)\|\cdot\|\eta\,\nabla_{\mathcal{L}}(\theta_{0};B)\|}$ (8)
which measures non-commutativity relative to step magnitude.
Random baseline control. To distinguish genuine geometric alignment from a trivial dimensionality artifact ($K$ dimensions out of $P$ will capture $\sim\!\sqrt{K/P}$ of any random vector), we compare $\rho_{\mathrm{exec}}$ against $\rho_{\mathrm{rand}}$: the projection fraction obtained from a random $K$-dimensional orthonormal basis (averaged over 10 trials). The ratio $\rho_{\mathrm{exec}}/\rho_{\mathrm{rand}}$ is our key diagnostic; values significantly above $1.0$ indicate that the execution subspace captures more commutator energy than expected by chance.
@@ 4.3 Localization of SGD Non-Integrability
[…]
Figure 5: Corrected commutator decomposition (PCA trajectory basis with random control). A: Projection fractions $\rho_{\mathrm{exec}}$ (green) and $\rho_{\mathrm{rand}}$ (red, dashed); the execution basis captures a small but significantly above-random fraction of commutator energy. B: Perpendicular component fraction $\|\delta_{\perp}\|/\|\delta\|$ (blue); $92\%$ throughout training, rising to $98\%$ late. C: Exec/random ratio $\rho_{\mathrm{exec}}/\rho_{\mathrm{rand}}$ (purple); peaks at $\sim\!14\times$ early, decays to $\sim\!2\times$ late, with the dashed line at $1.0$ marking the random expectation. D: Raw normalized defect $D$ (orange); grows throughout training, confirming that non-commutativity intensifies even after convergence.
SGD updates in the full high-dimensional parameter space exhibit strong non-commutativity. Figure 3 (bottom-right panel) shows the normalized commutator defect $D$ throughout training. Despite loss convergence by step $\sim\!5000$, the defect continues to grow with intermittent spikes, indicating that non-commutativity is not simply a manifestation of optimization difficulty but a persistent feature of the learning dynamics.
Projection reveals structure. To test whether this non-commutativity has geometric structure, we project commutators onto the learned execution subspace built from PCA of the weight trajectory (Section 3.3.1). Concretely, for each measurement step, we:
  1. 1.
    Construct the execution basis $B\in\mathbb{R}^{P\times K}$ from the top-$K$ PCA components of the accumulated weight trajectory, QR-orthonormalized;
  2. 2.
    Compute $\rho_{\mathrm{exec}}=\|BB^{T}\delta\|/\|\delta\|$, the fraction of the commutator $\delta=\theta_{AB}-\theta_{BA}$ captured by the execution subspace;
  3. 3.
    Compute $\rho_{\mathrm{rand}}$ by averaging the same projection fraction over $10$ random $K$-dimensional orthonormal bases.
Figure 4 shows the result: $\rho_{\mathrm{exec}}$ (green) is consistently small ($\approx 0.02$$0.13$) but systematically above $\rho_{\mathrm{rand}}$ (red, dashed), confirming that the vast majority ($92\%$) of non-commutativity lives in orthogonal staging directions, while the fraction within execution directions is small but geometrically structured.
Comparison against random baseline. The raw projection fraction alone does not distinguish genuine geometric alignment from a trivial dimensionality effect: any $K$-dimensional subspace of a $P$-dimensional space captures $\sim\!\sqrt{K/P}$ of a random vector’s norm. The critical diagnostic is the ratio $\rho_{\mathrm{exec}}/\rho_{\mathrm{rand}}$ (Figure 5, panel C). Early in training (first 20%), this ratio reaches $\mathbf{9.7\times}$ (median): the execution subspace captures nearly $10\times$ more commutator energy than a random subspace of equal dimension. Late in training (last 20%), the ratio decreases to $\mathbf{2.1\times}$, indicating that non-commutativity progressively leaves the execution manifold as the model converges.
This temporal pattern—commutators initially aligned with execution directions, then rotating away—has a natural interpretation. During early learning, gradient updates are shaped by the task structure and therefore align with the execution manifold. As the model converges, the residual non-commutativity becomes increasingly orthogonal, reflecting optimization interference in staging dimensions that does not affect the learned computation.
@@
Geometric role of overparameterization. The localization of non-integrability to orthogonal directions ($92\%$ perpendicular throughout, rising to $98\%$ late in training; Figure 5, panel B) has implications for the functional role of excess parameters. Extra dimensions provide a “buffer space” where optimization noise, task interference, and stochastic fluctuations can be sequestered without disrupting the core computational trajectory. The execution manifold acts as a stable attractor along which learning progresses, with the decreasing exec/random ratio indicating that residual non-commutativity is progressively expelled from execution directions as the model converges.
@@ 4.6 Extension to Modular Multiplication
Preliminary experiments with modular multiplication confirm that the execution manifold framework extends beyond addition. Multiplication tasks converge to slightly higher-dimensional manifolds (4466 dimensions versus 3344 for addition), with dimensionality scaling predictably with computational complexity. All attention matrices again operate in low-dimensional subspaces, and the same geometric principles apply: commutator localization to orthogonal directions, preferential alignment of non-commutativity with execution subspaces relative to random baselines, and curriculum effects. This suggests that manifold dimensionality may be a robust, task-dependent quantity that reflects inherent computational structure.
@@ 5.1 Theoretical Interpretation: Overparameterization and Parallel Exploration
The commutator analysis reveals a nuanced temporal picture: early in training, non-commutativity is preferentially aligned with the execution manifold ($\rho_{\mathrm{exec}}/\rho_{\mathrm{rand}}\approx 10\times$), reflecting task-structured gradient interference. As training converges, this ratio decreases to $\approx 2\times$, indicating that residual non-commutativity progressively leaves the execution subspace. Throughout training, the perpendicular component accounts for $92\%$ of total commutator magnitude. This localization provides a mechanism for stabilization: staging directions absorb the non-commutative aspects of SGD, allowing continued exploration without disrupting the core computational trajectory.
@@ 6 Conclusion
We have presented evidence that transformer learning dynamics collapse onto low-dimensional execution manifolds, with dimension 3344 for modular addition tasks despite ambient parameter space dimension $128$. This geometric structure unifies several empirical phenomena: attention bubbling emerges as saturation along routing coordinates, SGD commutators are preferentially aligned with the execution subspace early in training ($\rho_{\mathrm{exec}}/\rho_{\mathrm{rand}}\approx 10\times$) before rotating into orthogonal directions as training converges ($\approx 2\times$), and sparse features capture auxiliary routing structure distinct from distributed execution.
The localization of most ($92\%$) SGD non-commutativity to orthogonal staging directions, confirmed by comparison against random-subspace baselines, suggests a fundamental role for overparameterization: extra dimensions absorb optimization interference, allowing core computation to proceed via approximately path-independent updates. Training curriculum critically determines whether models discover unified execution manifolds supporting compositional generalization or become trapped in fragmented solution regions.